OpenStudy (anonymous):

what is the point- slope form of (-2,11) and (6,7)

OpenStudy (alexwee123):

well you would find the slope of (-2,11) and (6,7) first

OpenStudy (anonymous):

i might have it right i got y+4=-1/2(x-8)

OpenStudy (anonymous):

thanks

OpenStudy (alexwee123):

wait no

OpenStudy (alexwee123):

you plug in (-2,11) or (6,7)

OpenStudy (alexwee123):

not what you plugged in :/

OpenStudy (anonymous):

Slope is: \[m=\frac{y_1-y_2}{x_1-x_2}\] \[m=\frac{11-7}{-2-6}\] \[m=\frac{4}{-8}\] \[m=-0.5\] So you are right. Then for point-slope form; \[y-y_1=m(x-x_1)\] Then just plug in a set of points into y1 and x1, ex: \[y-7=-0.5(x-6)\]

OpenStudy (anonymous):

oh i had it backwards

OpenStudy (anonymous):

Yeah you can plug either point doesnt' matter

OpenStudy (anonymous):

got it